OpenStudy (anonymous):

The radius of a circle increases from 2m to 2.02 m. Use differential to approximate the change in area.

Parth (parthkohli):

\[A(r) = \pi r^2\]\[dA/dr = 2\pi r \]\[dA = 2\pi r dr \]

Parth (parthkohli):

\(dr\) and \(dA\) can be understood as small changes in radius and area respectively. With that sorted, are you given any of those two? And which one do you have to find?